4g of CaCl2 is dissolved in a 200cm^3 of a solution. Find the Molarity of CaCl2in the solution
1
Expert's answer
2020-07-24T14:52:24-0400

Molarity = moles/vol.of soln. in litre

In 4 gm. CaCl2

No. Of moles = 4/111 (mol.wt.of CaCl2is 111)

1Cm3 = 1ml.

Molarity of CaCl2 = (4*1000)/(111*200)

= 20/111

= 0.180 moles / litre